Determine the solution set of the inequality x^(2)-3x-10>0


1
Expert's answer
2021-02-03T02:33:58-0500

solution:

given; x2-3x-10 > 0

"\implies" x2-5x+2x-10 > 0

"\implies" x(x-5)+2(x-5) > 0

"\implies" (x+2)(x-5) > 0

case - "\Iota"

(x+2) > 0 & (x-5) > 0

x > -2 & x > 5

thus the common solution is x > 5

case - "\Iota" "\Iota"

(x+2) < 0 & (x-5) < 0

x < -2 & x < 5

thus the common solution is x < -2

hence,

using case - I and case - II we get x lies in ( -"\infty" ,-2) "\bigcup" (5,"\infty" )

solution set : { x | x"\in" ( -"\infty" ,-2) "\bigcup" (5,"\infty" ) }